# Lifelong Learning of Video Diffusion Models From a Single Video Stream

### v2

@@ Abstract
This work demonstrates that training autoregressive video diffusion models from a single, continuous video stream is not only possible but remarkably can also be competitive with standard offline training approaches given the same number of gradient steps. Our demonstration further reveals that this main result can be achieved using experience replay that only retains a subset of the preceding video stream. We also contribute three new single video generative modeling datasets suitable for evaluating lifelong video model learning: Lifelong Bouncing Balls, Lifelong 3D Maze, and Lifelong PLAICraft. Each dataset contains over a million consecutive frames from a synthetic environment of increasing complexity. [^1] Code and datasets will be released on acceptance.
@@ 1 Introduction
Modern AI systems do not learn in this manner but instead rely on a highly effective compromise: stochastic gradient descent from data streams made up of independently and identically distributed (i.i.d.) samples. Language models [71, 53], world models [33, 34] , and video models [35, 40, 5] are trained on random batches of short temporally-correlated segments, an approach that preserves short-range autocorrelations while approximating i.i.d.-ness through permutations. While effective, this approximately-i.i.d. learning paradigm does not offer a satisfying mechanism for updating models when new data arrive. Existing model updating techniques that are often used in practice – training from a checkpoint on the union of old and new data [3], fine-tuning on just the new data [80], or training completely anew from scratch [60] – do not operate in the training regime from which humans learn and suffer from problems such as high computation cost and forgetting [75].
@@
Alternatively, SGD on autocorrelated data streams is considered by some to be a viable candidate for human-like lifelong learning [49]. While there is a raft of work indicating that gradient-based learning on autocorrelated data is possible [28, 43, 57], folk wisdom maintains that this learning setup is hard, impractical, and prone to failure, especially for deep networks. Evidence of these beliefs can be found throughout the literature. The optimization community has developed numerous mechanisms for alleviating the effects of data stream temporal dependencies [46, 31, 19], suggesting the existence of problems with learning from autocorrelated data streams. Additionally, the Bayesian learning community has proposed numerous continual learning approaches through posterior updating [6, 13, 54, 7], though these theoretically sound approaches lack in practical scaled results. These bodies of work suggest a strong desire towards avoiding non-i.i.d. learning, making it reasonable to assume that lifelong learning of video models from a single autocorrelated video stream might be difficult if not impossible.
The primary contribution of this paper is simple yet profound. We empirically demonstrate that diffusion-based video models can be learned in a lifelong manner from a single autocorrelated video stream, obtaining comparable performance to standard (offline) i.i.d. training. To the best of our knowledge, this is the first demonstration of non-i.i.d. training of video diffusion models. Remarkably, successful lifelong learning does not require a complex setup. We find that a minimal set of established continual learning techniques, such as experience replay with limited memory, is sufficient to make lifelong- and i.i.d.-trained models indistinguishable in their performance given the same numbers of gradient steps and batch size.
(a) Lifelong Bouncing Balls frames subsampled from 5 second long videos.
(b) Lifelong 3D Maze frames subsampled from 2.5 second long videos.
(c) Lifelong PLAICraft frames subsampled from 4 second long videos.
@@
Figure 1: Ground truth video frames (1st row of each subfigure), offline learned models’ generated videos (2nd and 3rd rows), and lifelong learned models’ generated video frames (4th and 5th rows) for our datasets. The left two columns highlighted in red show a subset of frames used to condition the model. Videos generated by lifelong learned models trained with experience replay are diverse, visually plausible, and crucially indistinguishable from those of offline learned models in quality.
@@
As a secondary contribution, we introduce three novel lifelong learning video datasets with varying levels of temporal correlation, data repetitiveness, rare events, and non-stationarity. Even under an academic computational budget, we observe stable learning and reliable short-range extrapolations on all three benchmarks. As video modeling is a component of world modeling, our findings have the potential to open up new life-like approaches to learning, planning, and control in embodied AI agents.
@@ 2 Background
Video Diffusion Models (VDMs) are a class of generative models capable of synthesizing high-quality, temporally consistent videos [78, 40, 35, 41, 32, 10, 9, 14]. Rooted in the principles of denoising diffusion, video diffusion models progressively refine random noise into coherent video frames across a series of iterative denoising steps. Extending image diffusion models to videos requires capturing temporal dependencies between frames while preserving single-frame quality. These requirements present unique challenges, including the computational overhead of processing a large number of frames as well as developing architectures that can capture both spatial and temporal coherence efficiently. To handle long-duration videos, models learn the conditional probability of new frames given previously generated frames and then generate videos in an auto-regressive manner [35, 24].
@@
The goal of lifelong or continual learning is enabling models to continuously learn from new data with minimal forgetting of what was learned before [23, 79, 73, 81]. Approaches for lifelong learning include using regularization terms to penalize changes to parts of the network that encode previously learned information [45, 85, 48], improving the plasticity or the stability of the optimization algorithm [25, 38, 82], and enforcing the encoding of different tasks in mostly non-overlapping or orthogonal parts of the model [64, 66, 84]. Another popular approach is replay, whereby samples representative of past data are revisited by the model along with the current training data [61]. Such replayed samples can be obtained from a generative model [67], but often they are sampled from a memory buffer containing past training data, an approach referred to as experience replay [20, 62, 15, 2].
@@
Most work on neural network lifelong learning has focused on the highly simplified problem setting where a classification model is trained on a sequence of non-overlapping tasks, with each task seen only once. Often the model can train on each new task until convergence (referred to as the offline setting), although some works only allow a single pass over the data of each task (the online setting) [1, 22]. Beyond classification, lifelong learning research has explored the continual training of generative models, including generative adversarial networks, variational auto-encoders and diffusion models for image generation [86, 47, 29, 69]. However, as far as we are aware, no work thus far has explored lifelong learning of video diffusion models. The lifelong learning community has also shown interest in moving beyond learning in a strictly task-based manner. To create data streams with more complex temporal correlations, blurry task boundaries [4, 52] and repetition of previously seen concepts [37] have been used. Benchmarks for lifelong learning have also been constructed based on data from autonomous driving [74], using image datasets collected through time [11] and by concatenating thousands of short videos [18].
3 Datasets
Exploring the possibility of learning video models in a lifelong fashion requires long continuous video streams. As discussed in Section 2, existing video datasets usually consist of many short videos that may or may not be temporally related. To that end, we introduce three new video datasets, each of which is constructed from a single video stream without semantic discontinuities. The three datasets – Lifelong Bouncing Balls, Lifelong 3D Maze, and Lifelong PLAICraft video datasets – vary in terms of complexity, stochasticity, temporal correlation, and degree of non-stationarity.
3.1 Lifelong Bouncing Balls
We introduce two versions of the Lifelong Bouncing Balls dataset: Version O for “original” and C for “changing”.
[…]
Lifelong Bouncing Balls (C) introduces non-stationarity on top of Lifelong Bouncing Balls (O). While the balls’ motion matches that of (O), the blue channel values of all ball colors increase over time at a constant rate (red, yellow, and green ball colors respectively become fuchsia, white, and aqua by the end of the video). The evaluation set contains frames with balls of all previously observed colors. Solving Lifelong Bouncing Balls (C) requires learning the deterministic ball dynamics and color transitions from a correlated video stream where some details are unrepeated.
3.2 Lifelong 3D Maze
The Lifelong 3D Maze dataset contains 1 million 64x64 RGB video frames for training ($\sim$14 hours long at 20FPS) and 100,000 video frames for evaluation. The video is a first-person view of an agent that navigates a randomly generated 3D maze. Whenever the agent solves a maze, the walls of the solved maze come down and the walls of an unseen maze rise, at which point the agent attempts to solve the new maze. The mazes contain various sparsely appearing objects, including polyhedral gray rocks that flip the agent upside down upon being touched. Since the maze states are partially observable and their transition dynamics are stochastic, the future frames cannot be perfectly predicted given the past frames. Solving Lifelong 3D Maze requires learning the first-person sensory inputs associated with navigating a stochastically generated 3D environment and modeling infrequent events from a largely repetitive and correlated video stream.
3.3 Lifelong PLAICraft
The Lifelong PLAICraft dataset contains 1.85 million 1280x768 RGB video frames for training ($\sim$54 hours long at 10FPS) and 600,000 video frames for evaluation. To make training computationally feasible, we encode these video frames using the Stable Diffusion perceptual encoder [63] to a shape of 4x160x96. The training video is a first-person view of an anonymous player with an in-game ID of “Alex” engaged in the PLAICraft project’s multiplayer Minecraft survival world [36]. The evaluation video is a first-person view of another anonymous player with an in-game ID of “Kyrie” who explores the same Minecraft world. The videos capture multiple continuous play sessions spanning several months within this shared multiplayer survival world, showcasing various biomes in all three world dimensions, mining, crafting activities, construction, mob fighting, and player-to-player interactions. The Minecraft world contains aspects that repeat (ex. day-night cycle, players visiting their homes) and do not repeat (ex. felled trees, player chat logs). Solving Lifelong PLAICraft requires learning a highly nonstationary environment that changes in multiple timescales from a single correlated video stream.
@@ 4 Model and Training Regime
We use a U-Net video diffusion model with spatiotemporal attention introduced by Harvey et al. [35]. The video diffusion model’s U-Net operates on $K$ video frames at a time, where $K=10$ for Bouncing Balls and $K=20$ for the other two datasets. The first $K/2$ video frames are model inputs and the next $K/2$ video frames are prediction targets. During training, given the first $K/2$ uncorrupted video frames and the second $K/2$ video frames corrupted with Gaussian noise, the model regresses the values of the Gaussian noise [56]. During sampling, given the first $K/2$ clean video frames and the second $K/2$ video frames filled with Gaussian noise, the model iteratively “denoises” the second $K/2$ video frames until they resemble a plausible continuation of the first $K/2$ video frames.
@@
As a baseline representative of standard video models, we train models in an Offline Learning regime. At each training step, the model has access to $K$ contiguous video frames sampled from a random point in the data stream. The resulting training data resembles i.i.d. segments of $K$ video frames, even though all frames originate from a single autocorrelated stream. When using batch size $N_{\text{batch}}>1$ (as is common), this i.i.d. sampling is repeated $N_{\text{batch}}$ times.
@@
While the Offline Learning regime randomly permutes the data during training, the Lifelong Learning regime presents data to the model in order. This setup reflects the desiderata outlined in Section 1, where the model learns from a real-time data stream. At training step $t$, the model receives the $t$-th to $t\!+\!K$-th video frames of the (single) video that defines the dataset. After the model has taken a gradient step, the model receives the next set of frames ($t+1$ to $t\!+\!K\!+\!1$).

### v3

@@ Abstract
This work demonstrates that training autoregressive video diffusion models from a single video stream—resembling the experience of embodied agents—is not only possible, but can also be as effective as standard offline training given the same number of gradient steps. Our work further reveals that this main result can be achieved using experience replay methods that only retain a subset of the preceding video stream. To support training and evaluation in this setting, we introduce four new datasets for streaming lifelong generative video modeling: Lifelong Bouncing Balls, Lifelong 3D Maze, Lifelong Drive, and Lifelong PLAICraft, each consisting of one million consecutive frames from environments of increasing complexity. [^1] Code and datasets will be released on acceptance.
@@ 1 Introduction
Modern AI systems do not learn in this manner but instead rely on an effective compromise: stochastic gradient descent from data streams made up of independently and identically distributed (i.i.d.) samples. Language models [73, 56], world models [32, 33, 15, 1], and video models [35, 40, 7] are trained on random batches of short temporally correlated segments, an approach that preserves short-range autocorrelations while approximating i.i.d.-ness through permutations. While effective, this approximately i.i.d. learning paradigm does not offer a satisfying mechanism for updating models when new data arrive. Existing model updating techniques that are often used in practice—training from a checkpoint on the union of old and new data [5], fine-tuning on just the new data [84], or training completely anew from scratch [62] – do not operate in the training regime from which humans learn and suffer from problems such as high computation cost and forgetting [78].
@@
Alternatively, SGD on autocorrelated data streams is considered by some to be a viable candidate for human-like lifelong learning [49]. While there is a raft of work indicating that gradient-based learning on autocorrelated data is possible [27, 43, 59], folk wisdom maintains that this learning setup is hard, impractical, and prone to failure, especially for deep networks. Evidence of these beliefs can be found throughout the literature. The optimization community has developed numerous mechanisms for alleviating the effects of data stream temporal dependencies [46, 30, 19], suggesting the existence of problems with learning from autocorrelated data streams. The Bayesian learning community has proposed numerous continual learning approaches through posterior updating [8, 14, 57, 9], though these approaches lack practical scaled results. These bodies of work suggest a strong desire to avoid non-i.i.d. learning, making it reasonable to assume that lifelong learning of video models from a single autocorrelated video stream might be highly challenging.
Refer to caption (a) Lifelong Bouncing Balls videos. Refer to caption (b) Lifelong 3D Maze videos. Ground Truth Offline Learning Sample 1 Offline Learning Sample 2 Lifelong Learning Sample 1 Lifelong Learning Sample 2
Refer to caption (c) Lifelong Drive videos. Refer to caption (d) Lifelong PLAICraft videos. Ground Truth Offline Learning Sample 1 Offline Learning Sample 2 Lifelong Learning Sample 1 Lifelong Learning Sample 2
@@
Figure 1: Ground truth video frames (1st row of each subfigure), offline learned models’ generated videos (2nd and 3rd rows), and lifelong learned models’ generated video frames (4th and 5th rows) for our datasets. The leftmost columns highlighted in red show the frame conditioned upon the model. Videos generated by lifelong learned models trained with experience replay are diverse, visually plausible, and are comparable to offline models in visual fidelity. Figures are best viewed zoomed in.
The primary contribution of this paper is an empirical proof-of-concept that diffusion-based video models can learn in a lifelong manner from a single autocorrelated video stream while obtaining similar performance to standard (offline) i.i.d. training. We demonstrate successful lifelong learning even on highly nonstationary partially-observable 3D domains across multiple datasets, model architectures, parameter sizes, and optimizers. Remarkably, successful lifelong learning does not require a complex setup. We find that a minimal set of established continual learning techniques, such as experience replay with limited memory, is sufficient to make lifelong- and i.i.d.-trained models qualitatively indistinguishable given the same numbers of gradient steps and batch size.
@@
As a secondary contribution, we introduce four single video lifelong learning datasets with varying levels of temporal correlation, repetitiveness, perceptual complexity, and non-stationarity. Even with an academic computational budget, we observe stable learning and reliable short-range extrapolations on every dataset. As video modeling is a component of world modeling, our findings have the potential to open up new life-like streaming approaches to learning, planning, and control in embodied agents.
@@ 2 Background
Video Diffusion Models (VDMs) are a class of generative models capable of synthesizing high-quality, temporally consistent videos [82, 40, 35, 41, 31, 11, 10, 52, 15]. Rooted in the principles of denoising diffusion, video diffusion models progressively refine random noise into coherent video frames across a series of iterative denoising steps. Extending image diffusion models to videos requires capturing temporal dependencies between frames while preserving single-frame quality, which existing work typically achieves using attention mechanisms.
@@
The goal of lifelong or continual learning is enabling models to continuously learn from new data with minimal forgetting of what was learned before [23, 83, 76, 85]. Approaches to lifelong learning include using regularization to penalize changes to parts of the network that encode previously learned information [45, 89, 48], improving the plasticity or stability of the optimization algorithm [24, 38, 86], and enforcing the encoding of different tasks in orthogonal parts of the model [66, 68, 88]. Another popular approach is replay, whereby the model revisits samples representative of past data along with the current training data [63]. Such replayed samples can be obtained from generative models [69], but often they are sampled from a memory buffer containing past training data, an approach referred to as experience replay [20, 64, 16, 4].
Most work on neural network lifelong learning has focused on the highly simplified problem setting where a classification model is trained on a sequence of non-overlapping tasks, with each task seen only once. Often the model can train on each new task until convergence (referred to as the offline setting), although some works only allow a single pass over the data of each task (the online setting) [2, 22]. Beyond classification, lifelong learning research has explored the continual training of generative models, including GANs, VAEs and diffusion models for image generation [90, 47, 28, 71]. However, no work thus far has explored lifelong learning of video diffusion models. The lifelong learning community has also shown interest in moving beyond learning in a strictly task-based manner. To create data streams with more complex temporal correlations, blurry task boundaries [6, 55] and repetition of previously seen concepts [37] have been used. Benchmarks for lifelong learning have also been constructed based on data from autonomous driving [77], using image datasets collected through time [12] and by concatenating thousands of short videos [18].
3 Datasets
Exploring the possibility of learning video models in a lifelong fashion requires (1) long continuous video streams and (2) varying levels of complexity to gauge challenges and limitations. As discussed in Section 2, existing video datasets usually consist of many short videos that may or may not be temporally related. To that end, we introduce four new video datasets derived from single video streams without semantic discontinuities that vary in perceptual complexity, stochasticity, temporal correlation, and degree of non-stationarity (refer to Appendix C for more details).
Lifelong Bouncing Balls
We present two versions of Lifelong Bouncing Balls: Versions O and C.
Lifelong Bouncing Balls (O) contains 1 million 32x32 RGB video frames for training ($\sim$28 hours long at 10FPS) and another 1 million video frames for evaluation. The video contains two colored balls that deterministically bounce around in a 2D environment, colliding with boundaries and each other. Upon collision, each ball changes velocity according to the conservation of momentum and cycles through a repeating color sequence of red, yellow, red, and green. Since the ball states are fully observable and the transition dynamics are deterministic, the future frames can be perfectly predicted. Solving Lifelong Bouncing Balls (O) requires learning the 2D environment’s deterministic dynamics and retaining the balls’ color transition histories from a correlated and repetitive video stream.
Lifelong Bouncing Balls (C) introduces non-stationarity on top of Lifelong Bouncing Balls (O). While the balls’ motion matches that of (O), the blue channel values of all ball colors increase over time at a constant rate (red, yellow, and green ball colors respectively become fuchsia, white, and aqua by the end of the video). The evaluation set contains frames with balls of all previously observed colors. Solving Lifelong Bouncing Balls (C) requires learning the deterministic ball dynamics and color transitions from a correlated video stream with unrepetitive details.
Lifelong 3D Maze
The Lifelong 3D Maze dataset contains 1 million 64x64 RGB video frames for training ($\sim$14 hours long at 20FPS) and 100,000 video frames for evaluation. The video is a first-person view of an agent that navigates a randomly generated 3D maze. Whenever the agent solves a maze, the walls of the solved maze come down and the walls of an unseen maze rise, at which point the agent attempts to solve the new maze. The mazes contain various sparsely appearing objects, including polyhedral gray rocks that flip the agent upside down upon being touched. Since the maze states are partially observable and their transition dynamics are stochastic, the future frames cannot be perfectly predicted given the past frames. Solving Lifelong 3D Maze requires learning the first-person sensory inputs associated with navigating a stochastic 3D environment and modeling infrequent events from a largely repetitive and correlated video stream.
Lifelong Drive
The Lifelong Drive dataset contains 1 million 512x512 RGB video frames for training ($\sim$14 hours long at 20FPS) and 100,000 video frames for evaluation. We encode these video frames using the Stable Diffusion perceptual encoder [65] to a shape of 4x64x64. The video is continuous dashcam footage of a vehicle traveling on a series of highways from Chongqing to Shanghai for 1675 kilometers [79]. The drive contains a diverse range of scenery such as mountains, cities, and tunnels, weather conditions such as sunny, cloudy, and rainy weather, and other cars and trucks moving at different speeds. Solving Lifelong Drive requires modeling real world perceptual data from a single correlated and nonstationary video stream.
Lifelong PLAICraft
The Lifelong PLAICraft dataset contains 1 million 1280x768 RGB video frames for training ($\sim$28 hours long at 10FPS) and 500,000 video frames for evaluation. We encode these video frames using the Stable Diffusion perceptual encoder to a shape of 4x160x96. The training video is a first-person view of an anonymous player with an in-game ID of “Alex” engaged in the PLAICraft project’s multiplayer Minecraft survival world [36]. The evaluation video is a first-person view of another anonymous player with an in-game ID of “Kyrie” who explores the same world. The videos capture multiple continuous play sessions spanning several months within this shared multiplayer survival world, showcasing various biomes in all three world dimensions, mining, crafting activities, construction, mob fighting, and player-to-player interactions. The Minecraft world contains aspects that repeat (ex. day-night cycle, players visiting their homes) and do not repeat (ex. felled trees, player chat logs). Solving Lifelong PLAICraft requires learning a highly nonstationary environment that changes in multiple timescales from a single correlated video stream.
@@ 4 Model and Training Regime
We use video diffusion models with U-Net [35] and Transformer [52] neural architectures. The diffusion models’ networks operate on $K$ video frames at a time, where $K=10$ for Lifelong Bouncing Balls and Lifelong PLAICraft while $K=20$ for the Lifelong 3D Maze and Lifelong Drive. The first $K/2$ video frames are model inputs and the next $K/2$ video frames are prediction targets.
During training, given the first $K/2$ uncorrupted video frames and the second $K/2$ video frames corrupted with Gaussian noise, the model regresses the values of the Gaussian noise [39]. Specifically, let $\boldsymbol{x}$ be a size $K$ window of consecutive video frames, and $F_{\boldsymbol{\theta}}$ be the denoising neural network. The denoising loss function is defined as
$\ell(\boldsymbol{\theta},\boldsymbol{x})=\mathbb{E}_{\boldsymbol{\epsilon},s}\left\|\boldsymbol{\epsilon}-F_{\boldsymbol{\theta}}(\boldsymbol{x}^{\mathrm{obs}},\boldsymbol{x}^{\mathrm{lat}}_{s},s)\right\|^{2},$ (1)
where $\epsilon\sim\mathcal{N}(0,I)$ is a unit normal random vector, and $s\sim\mathcal{U}(1,S)$ is the diffusion noising timestep. The superscripts $\mathrm{obs}$ and $\mathrm{lat}$ respectively represent the input and predicted part of $\boldsymbol{x}$ i.e., the first and second $K/2$ frames. The subscript $s$ denotes that the video frames are corrupted to the $s$-th diffusion timestep using $\epsilon$ as noise [39]. During sampling, given the first $K/2$ clean video frames $\boldsymbol{x}^{\mathrm{obs}}$ and the second $K/2$ video frames $\boldsymbol{x}^{\mathrm{lat}}$ filled with Gaussian noise, the model iteratively denoises $\boldsymbol{x}^{\mathrm{lat}}$ to produce a plausible continuation of $\boldsymbol{x}^{\mathrm{obs}}$ using Karras et al. [44]’s stochastic sampler.
For the rest of this section, we represent the entire video stream as $\mathcal{X}$ and the window of $K$ video frames starting from the $i^{\mathrm{th}}$ frame as $\mathcal{X}^{i:i+K}$. In addition, the expectation in eq. 1 is approximated with a single-sample Monte Carlo estimate.
@@
As a baseline representative of standard video models, we train models in an Offline Learning regime. The $t$-th training step loss is $\mathcal{L}^{\mathrm{offline}}_{t}(\boldsymbol{\theta},\mathcal{X})=\ell(\boldsymbol{\theta},\mathcal{X}^{i:i+K})$, where $i$ is a uniformly sampled video frame index. The resulting training data resembles i.i.d. segments of $K$ video frames, even though all frames originate from a single autocorrelated stream. When using batch size $N>1$ (as is common), this i.i.d. sampling is repeated $N$ times.
@@
While the Offline Learning regime randomly shuffles the data during training, the Lifelong Learning regime presents data to the model in order. This setup reflects the desiderata from Section 1, where the model learns from a real-time data stream. The $t$-th training step loss is $\mathcal{L}^{\mathrm{lifelong}}_{t}(\boldsymbol{\theta},\mathcal{X})=\ell(\boldsymbol{\theta},\mathcal{X}^{t:t+K})$, i.e., the model receives a sliding window of frames from the single video $\mathcal{X}$. When using batch size $N>1$, we fill the batch with $N-1$ copies of the current video $X^{t:t+K}$. This results in a lower variance estimate of eq. 1 by sampling multiple values of $\epsilon$ and $s$.
This Lifelong Learning setup is compatible with numerous techniques proposed by the continual learning community. As a demonstration, we augment the online stream with a replay buffer [20] that retains a buffer $\mathcal{M}$ of past video subsequences chosen through reservoir sampling [81], where the buffer size is chosen to be a small fraction of the total data stream size. With a batch size of $N$, the new $t$-th training step loss is
$\mathcal{L}^{\mathrm{lifelong}}_{t}(\boldsymbol{\theta},\mathcal{X})=\frac{1}{N}\!\!\left(\!\ell(\boldsymbol{\theta},\mathcal{X}^{t:t+K}\!)+\!\!\sum_{i=1}^{N-1}\!\ell(\boldsymbol{\theta},\mathcal{M}^{i})\!\right)\!,$ (2)
where $\mathcal{M}^{i}$ is a randomly chosen video from the replay buffer. We refer to this algorithm as Lifelong Learning for the rest of this paper. We emphasize that replay is a simple technique that could be replaced with other continual learning methods.
5 Experiments
This section qualitatively and quantitatively compares offline learned and lifelong learned video diffusion models. The learning algorithms use the same batch size and number of gradient steps, equalizing the amount of computation and runtime memory. We refer the readers to Appendix B for additional experiment details. Video samples are available in the supplementary materials.
5.1 Lifelong Bouncing Balls
As discussed in Section 3, the Lifelong Bouncing Balls datasets feature repetitive and deterministic dynamics, with stationary (Version O) and non-stationary (Version C) color transitions.
Setup
The U-Net diffusion models for Lifelong Bouncing Balls (O) and Lifelong Bouncing Balls (C) respectively have 8 and 74 million parameters, whereas the Transformer diffusion models for both datasets have 42 million parameters. All models are trained with a batch size of 2 and are evaluated using 45 frames that they autoregressively generate conditioned on 5 ground truth frames. Lifelong Learning retains 5 percent of the stream’s frames in the replay buffer. We measure the sampled videos’ perceptual and temporal coherence using FVD [74] and report the loss to measure learning progress. We also measure the models’ understanding of the ball movement using minADE [61] and of the color transition using our novel ColorKL metric (Appendix A), extracting the ball positions and colors from the frames using 2D convolutions.
(a) Lifelong Bouncing Balls (C) videos.
(b) Lifelong 3D Maze videos.
(c) Lifelong Drive videos.
(d) Lifelong PLAICraft videos.
Figure 3: Qualitative result comparisons. The top, middle, and bottom rows depict videos from the ground truth data, offline learned model, and lifelong learned model respectively. Additional video samples are presented in the supplementary materials.